You are here: Foswiki>Commissioning Web>Logs>28Jan013_Log (29 Jan 2013, JuanGuerra)Edit Attach

UT 20130128 AO#1 AO#2

Telescope Operator: Steve..(LBTO)
SW Support: Cox,D (Tucson)
AO Support: JCGuerra(LBTO),JChristou(LBTO)
Telescope Support: Jeff U(LBTO)
Instrument Support: Tim S (LBTO)


AGW#2 in the telescope and AGW#1 in the lab. Almost done with the AGW#1 in the lab. Starting operation at 3pm. bad weather impossible to open. Night observations cancelled. Enrico Pinna injured his Achiles tendon and it was took down to Tucson during the night. The Arcetri team left this evening. Enrico will take it to UMC.


The AGW#2 in the telescope working with operations of the AO system and adding some notes to the AO user document for daytime. Revision and update the list of spares and devices that they have to be replaced. Updated the spare list.

Julian and Juan CArlos used the Wavefront sensor system and playing with different configurations based in star magnitude, the system responds very well for all the test and it showed one failure of the ccd47 that it was recovered changing the bin of the ccd47(bcu47 is reseted).

Juan CArlos performed a test leaving the WFS at close loop for about 30 minutes and we didn't see any failure.

-- JuanGuerra - 28 Jan 2013
Topic revision: r2 - 29 Jan 2013, JuanGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback